the angle between the pair of line s whose equation is axsquare+10xy+mysquare+++++5x+10y=0
 

Grade:12

1 Answers

Vikas TU
6 years ago
x2+10xy+y2+5x+10y=0 
Point between two straight lines is given by tan A =(m1-m2/1+m1m2) 
M1= slant of first line,m2 =slope of second line 
Or, then again we can specifically discover utilizing this technique. 
general condition of combine of staight line is ax2 + 2hxy + by2 + 2gx + 2fy + c = 0 
thus, a = 1, b = 1, h = 5, g = 5/2, f = 5, c = 0. 
to get edge, put require an incentive in the condition 
point = tan - 1 [ | 2 ( h2 – stomach muscle )1/2 |/| a + b | ] 
the appropriate response will be tan-1 [( 25 – )1/2/2] 
=tan-1√6 
x2+10xy+y2+5x+10y=0 
Point between two straight lines is given by tan A =(m1-m2/1+m1m2) 
M1= slant of first line,m2 =slope of second line 
Or, then again we can specifically discover utilizing this technique. 
general condition of combine of staight line is ax2 + 2hxy + by2 + 2gx + 2fy + c = 0 
thus, a = 1, b = 1, h = 5, g = 5/2, f = 5, c = 0. 
to get edge, put require an incentive in the condition 
point = tan - 1 [ | 2 ( h2 – stomach muscle )1/2 |/| a + b | ] 
the appropriate response will be tan-1 [( 25 – )1/2/2] 
=tan-1√6